# DiarizationLM: Speaker Diarization Post-Processing with Large Language Models

### v8

@@ Abstract
In this paper, we introduce DiarizationLM, a framework to leverage large language models (LLM) to post-process the outputs from a speaker diarization system. Various goals can be achieved with the proposed framework, such as improving the readability of the diarized transcript, or reducing the word diarization error rate (WDER). In this framework, the outputs of the automatic speech recognition (ASR) and speaker diarization systems are represented as a compact textual format, which is included in the prompt to an optionally finetuned LLM. The outputs of the LLM can be used as the refined diarization results with the desired enhancement. As a post-processing step, this framework can be easily applied to any off-the-shelf ASR and speaker diarization systems without retraining existing components. Our experiments show that a finetuned PaLM 2-S model can reduce the WDER by rel. $55.5\%$ on the Fisher telephone conversation dataset, and rel. $44.9\%$ on the Callhome English dataset [^1] This project is currently a work in progress. We may add new experiment results to this paper soon..
@@
Model: https://huggingface.co/google/DiarizationLM-13b-Fisher-v1
@@ 4.4 Zero-shot and one-shot baselines
Prompt prefix for zero-shot:
In the speaker diarization transcript below, some words are potentially misplaced.
Please correct those words and move them to the right speaker.
@@
Prompt prefix for one-shot:
In the speaker diarization transcript below, some words are potentially misplaced. Please correct those words and move them to the right speaker. For example, given this input transcript,
<spk:1> How are you doing today? I <spk:2> am doing very well. How was everything at the <spk:1> party? Oh, the party? It was awesome. We had lots of fun. Good <spk:2> to hear!
The correct output transcript should be:
<spk:1> How are you doing today? <spk:2> I am doing very well. How was everything at the party? <spk:1> Oh, the party? It was awesome. We had lots of fun. <spk:2> Good to hear!
@@ 4.6 Case studies
Table 1: Evaluation results of the USM + turn-to-diarize baseline system and the results post-processed by DiarizationLM. For DiarizationLM, we experiment with PaLM 2 foundation models with and without finetuning on the diarization task. We also experiment with Llama 2 and Llama 3 models finetuned on the diarization task. WERs are the same for all systems due to TPST. All numbers are percentages.
@@
+ PaLM 2-S finetuned (hyp2ora flavor) - 2.37 16.93 - 4.25 20.22
+ PaLM 2-S finetuned (deg2ref flavor) - 3.94 18.55 - 5.33 21.47
+ PaLM 2-S finetuned (mixed flavor) - 2.41 16.94 - 4.76 20.84
+ Llama 2 13B finetuned (mixed flavor) - 3.65 18.92 - 8.02 25.11
+ Llama 3 8B finetuned (mixed flavor) - 4.40 19.76 - 12.27 30.80
@@ Appendix A Open source models
Apart from PaLM 2 [46] models, we have also experimented with open source models such as Llama 2 [48] and Llama 3. These finetuned models are publicly available on Hugging Face.
@@
The evaluation results of this model is also provided in Table 1. We can see that similar to the finetuned PaLM 2-S models, this model also significantly reduces the WDER and cpWER on the Fisher testing set. However, on the Callhome testing set, we are seeing slight degradation in WDER and cpWER, indicating potential overfitting due to training for $\sim$4 epochs.
@@
The evaluation results of this model is also provided in Table 1. We can see that its performance is significantly worse than the Llama 2 13B model from Appendix A.1, but still better than the USM + turn-to-diarize baseline on Fisher testing set.

### v11

@@ Abstract
In this paper, we introduce DiarizationLM, a framework to leverage large language models (LLM) to post-process the outputs from a speaker diarization system. Various goals can be achieved with the proposed framework, such as improving the readability of the diarized transcript, or reducing the word diarization error rate (WDER). In this framework, the outputs of the automatic speech recognition (ASR) and speaker diarization systems are represented as a compact textual format, which is included in the prompt to an optionally finetuned LLM. The outputs of the LLM can be used as the refined diarization results with the desired enhancement. As a post-processing step, this framework can be easily applied to any off-the-shelf ASR and speaker diarization systems without retraining existing components. Our experiments show that a finetuned PaLM 2-S model can reduce the WDER by rel. $55.5\%$ on the Fisher telephone conversation dataset, and rel. $44.9\%$ on the Callhome English dataset.
@@
Model: https://huggingface.co/google/DiarizationLM-8b-Fisher-v2
Demo: https://huggingface.co/spaces/diarizers-community/DiarizationLM-GGUF
@@ 4.6 Case studies
Table 1: Evaluation results of the USM + turn-to-diarize baseline system and the results post-processed by DiarizationLM. For DiarizationLM, we experiment with PaLM 2 foundation models with and without finetuning on the diarization task. We also experiment with Llama 2 and Llama 3 models finetuned on the diarization task (model details in Appendix A). WERs are the same for all systems due to TPST. All numbers are percentages.
@@
+ PaLM 2-S finetuned (hyp2ora) - 2.37 16.93 - 4.25 20.22
+ PaLM 2-S finetuned (deg2ref) - 3.94 18.55 - 5.33 21.47
+ PaLM 2-S finetuned (mixed) - 2.41 16.94 - 4.76 20.84
+ Llama 2 13B finetuned (mixed) v1 - 3.65 18.92 - 8.02 25.11
+ Llama 3 8B finetuned (mixed) v1 - 4.40 19.76 - 12.27 30.80
+ Llama 3 8B finetuned (mixed) v2 - 3.28 18.37 - 6.66 23.57
@@ Appendix A Open source models
Apart from PaLM 2 [46] models, we have also experimented with open source models such as Llama 2 [48] and Llama 3. These finetuned models are publicly available on Hugging Face. The finetuning scripts are available at: https://github.com/google/speaker-id/tree/master/DiarizationLM/unsloth.
@@
The evaluation results of this model is also provided in Table 1. We can see that similar to the finetuned PaLM 2-S models, this model also significantly reduces the WDER and cpWER on the Fisher testing set. However, on the Callhome testing set, we are seeing degradation in WDER and cpWER, indicating overfitting on the Fisher dataset. We have found this is because the loss is computed on both the prompt and the completion tokens during training.
@@
The evaluation results of this model is also provided in Table 1. We can see that its performance is significantly worse than the Llama 2 13B model from Appendix A.1, but still better than the USM + turn-to-diarize baseline on Fisher testing set. Similarly, the performance on Callhome is worse than baseline, because the loss is computed on both the prompt and the completion tokens during training.
A.3 google/DiarizationLM-8b-Fisher-v2
This model is finetuned on the training subset of the Fisher corpus, using a LoRA adapter [63] of rank 256. The total number of training parameters is 671,088,640. With a batch size of 16, this model has been trained for 28800 steps, which is $\sim$9 epochs of the training data.
Note that when training this model, the loss is only computed on the completion tokens. This is the biggest different between this model and the previous model in Appendix A.2 ( google/DiarizationLM-8b-Fisher-v1).
The evaluation results of this model is also provided in Table 1. We can see that its performance is significantly better than the model from Appendix A.2. Specifically, we no longer see degradation on Callhome testin set.